OpenStudy (anonymous):

Find the solution of four times the square root of the quantity of x plus 2 equals negative 16

OpenStudy (anonymous):

\[4\sqrt{x+2} = -16\]

OpenStudy (triciaal):

first divide each side by 4 next square each side

OpenStudy (triciaal):

remember when you square a negative it becomes a positive

OpenStudy (texaschic101):

4 sq rt (x + 2) = -16 --- divide both sides by 4 sq rt (x + 2) = - 4 x + 2 = -4^2 x + 2 = 16 x = 16 - 2 x = 14......but I am not sure if it is (+/-) 14 or just 14

OpenStudy (anonymous):

\[\sqrt{x+2} =-4?\]

OpenStudy (anonymous):

oops, looks like you already posted :o

OpenStudy (triciaal):

it would just be 14 you only have +/- when taking the square root

OpenStudy (anonymous):

would 14 be extraneous or non-extraneous?

OpenStudy (texaschic101):

thanks for clearing that up triciaal :)

OpenStudy (anonymous):

alright so what I understand is that you have to plug it in, but I'm not sure of how I'd be doing the equation? I would do it pretty much the same way I'd be finding x?

OpenStudy (anonymous):

I found it to be extraneous c: may I ask another question please? >.<

OpenStudy (triciaal):

based on your ? using a different approach yes extraneous

OpenStudy (triciaal):

sure

OpenStudy (anonymous):

alright, thank you c: my question is how would I simplify this: \[9\div 3 - 2i\] it may seem easy but I'm rather lost with these

OpenStudy (triciaal):